Calculate the mass of NaH2PO4 (FW= 119.98) and of Na2HPO4 (FW= 141.96) needed to prepare 500 mL of 1.0M NaH2PO4/Na2HPO4 buffer at pH = 8.0.



1
Expert's answer
2021-10-30T11:16:36-0400

500ml/119.98=4.17mol

= 119.98+141.96= 261.94

= 119.98/261.94×4.17

1.91×119.98=229.17g

=141.96/261.94×4.17×141.96=320.82g